OpenStudy (anonymous):

Define ‘slope’. A. The slope is the change of y over the change of x. B. The slope is the change of x over the change of y. C. The slope is the y coordinate of a point over the x coordinate of the same point. D. The slope is the x coordinate of a point over the y coordinate of the same point.

OpenStudy (anonymous):

A is correct

OpenStudy (anonymous):

k i know what a slope is but having trouble choosing between the options at top

OpenStudy (anonymous):

thought so

OpenStudy (ash2326):

Slope is the change of y over the change of x \[Slope=\frac{y2-y1}{x2-x1}\]

OpenStudy (anonymous):

appreciate it, ty

Parth (parthkohli):

A formal definition is \[ Slope = {\Delta y \over \Delta x}\]

OpenStudy (anonymous):

\[\large Slope=\frac{y_2-y_1}{x_2-x_1} = \frac{\Delta {y}}{\Delta {x}}\]